# Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo

### v2

@@ Abstract
Quantum entanglement can reveal the fundamental organizing principles of quantum matter, ranging from non-Fermi liquids in high-temperature superconductors to black holes [1, 2]. However, obtaining the entanglement structure of realistic systems is a challenging task. When dealing with physically relevant situations that include finite temperature or the study of microscopic subregions, one has to work with mixed states (density matrices), and simple quantities like the von Neumann or Rényi entropies cannot be used as measures of entanglement. In the simplest case of bipartite entanglement, one can for example exploit the partial transpose to construct the (logarithmic) negativity [3, 4, 5], which is a true entanglement measure not polluted by classical correlations. However, it is challenging to compute in many-body systems for subregions with many degrees of freedom (including quantum field theories) where one can only obtain related quantities via the replica trick [6, 7]. These Rényi analogs are not entanglement measures and thus cannot be used to obtain the entanglement structure. A striking example is in boson/spin systems poised near a quantum critical point (QCP), where the logarithmic negativity (LN) decays faster than any power, as shown for CFTs in 1d [7] and higher $d$ [8], while the Rényi moments decay algrebraically. These are special cases of a general phenomenon regarding the “fate of entanglement” [9] where any multi-partite form of entanglement (including but more general than bipartite) entirely disappear as a system heats up, evolves in time, or as its parts grow separated.
@@ Acknowledgments
$\begin{pmatrix}0.22268(4)&0.13787(4)&0.13182(3)&0.10746(2)&0.13182(3)&0.10746(2)&0.11605(3)&0.12789(5)\\ 0.13787(4)&0.09911(4)&0.08899(3)&0.08646(2)&0.08899(3)&0.08646(2)&0.08746(4)&0.11605(3)\\ 0.13182(3)&0.08899(3)&0.08911(2)&0.08046(3)&0.08408(3)&0.07709(3)&0.08646(2)&0.10746(2)\\ 0.10746(2)&0.08646(2)&0.08046(3)&0.08911(2)&0.07709(3)&0.08408(3)&0.08899(3)&0.13182(3)\\ 0.13182(3)&0.08899(3)&0.08408(3)&0.07709(3)&0.08911(2)&0.08046(3)&0.08646(2)&0.10746(2)\\ 0.10746(2)&0.08646(2)&0.07709(3)&0.08408(3)&0.08046(3)&0.08911(2)&0.08899(3)&0.13182(3)\\ 0.11605(3)&0.08746(4)&0.08646(2)&0.08899(3)&0.08646(2)&0.08899(3)&0.09911(4)&0.13787(4)\\ 0.12789(5)&0.11605(3)&0.10746(2)&0.13182(3)&0.10746(2)&0.13182(3)&0.13787(4)&0.22268(4)\end{pmatrix}$

### v3

@@ Abstract
Quantum entanglement can reveal the fundamental organizing principles of quantum matter, ranging from non-Fermi liquids in high-temperature superconductors to black holes [1, 2]. However, obtaining the entanglement structure of realistic systems is a challenging task. When dealing with physically relevant situations that include finite temperature or the study of microscopic subregions, one has to work with mixed states (density matrices), and simple quantities like the von Neumann or Rényi entropies cannot be used as measures of entanglement. In the simplest case of bipartite entanglement, one can for example exploit the partial transpose to construct the (logarithmic) negativity [3, 4, 5, 6], which is a true entanglement measure not polluted by classical correlations. However, it is challenging to compute in many-body systems for subregions with many degrees of freedom (including quantum field theories) where one can only obtain related quantities via the replica trick [7, 8]. These Rényi analogs are not entanglement measures and thus cannot be used to obtain the entanglement structure. A striking example is in boson/spin systems poised near a quantum critical point (QCP), where the logarithmic negativity (LN) decays faster than any power, as shown for CFTs in 1d [8] and higher $d$ [9], while the Rényi moments decay algrebraically. These are special cases of a general phenomenon regarding the “fate of entanglement” [10] where any multi-partite form of entanglement (including but more general than bipartite) entirely disappear as a system heats up, evolves in time, or as its parts grow separated.
@@ Acknowledgments
$\begin{pmatrix}0.22268(5)&0.13781(3)&0.13781(3)&0.09887(3)&0.12589(4)&0.11605(3)&0.11605(3)&0.12789(5)\\ 0.13781(3)&0.09914(2)&0.09388(4)&0.08048(2)&0.08409(2)&0.09241(2)&0.08746(3)&0.11605(3)\\ 0.13781(3)&0.09388(4)&0.09914(2)&0.08048(2)&0.08409(2)&0.08746(3)&0.09241(2)&0.11605(3)\\ 0.09887(3)&0.08048(2)&0.08048(2)&0.07904(3)&0.06672(3)&0.08409(2)&0.08409(2)&0.12589(4)\\ 0.12589(4)&0.08409(2)&0.08409(2)&0.06672(3)&0.07904(3)&0.08048(2)&0.08048(2)&0.09887(3)\\ 0.11605(3)&0.09241(2)&0.08746(3)&0.08409(2)&0.08048(2)&0.09914(2)&0.09388(4)&0.13781(3)\\ 0.11605(3)&0.08746(3)&0.09241(2)&0.08409(2)&0.08048(2)&0.09388(4)&0.09914(2)&0.13781(3)\\ 0.12789(5)&0.11605(3)&0.11605(3)&0.12589(4)&0.09887(3)&0.13781(3)&0.13781(3)&0.22268(5)\end{pmatrix}$